Part II: SPSS Application

These questions require the use of SPSS. Remember you must submit all of your work within this word document. You will need to take a screen shot of your data view if necessary, or copy and paste your output into the spaces below. Remember to report the exact p value provided by SPSS output – simply reporting p<.05 or p>.05 is not acceptable (unless SPSS output states p=.000 – in that case you can report p<.001).

Research Scenario: A developmental psychologist wants to study the effect of instructional type on recall of the alphabet. She randomly selects a sample of 20 students at the local preschool and assigns them to two groups, an “Individual” group who recites and practices the alphabet on their own using personal learning devices (n=10) and a “Group” group who studies the alphabet together (n=10). After one week of daily 30 minute sessions, all of the children are asked to recite the alphabet. The number of correctly recalled letters for each group are presented below. Create your SPSS data file (remember that there will be a “grouping variable” for instructional type and your second variable, or column, entered into SPSS will be your dependent variable, recall. Look at the example file from this week’s presentation if you want a visual reminder). Conduct an independent-samples t test to determine whether the instructional types resulted in a difference in recall of the alphabet. (14 pts)

Individual

Group

20

13

11

14

16

20

9

14

22

12

23

14

19

9

18

21

8

11

12

15

15)

Paste the appropriate SPSS output. (3 pts)

Answer:

16)

Paste appropriate SPSS graph.    (3 pts)

Answer:

17)

Calculate the effect size given the SPSS output as shown in this week’s presentation. Show all of your work. (2 pts)

Work and Answer:

18)

Write an APA-style Results section based on your analyses. Remember to use complete sentences, include the statistical notation, and include a decision about the null hypothesis. If the results are significant, include the effect size and its interpretation as well. (4 pts)

Answer:

19)

If the null hypothesis had been directional, what significance value would you report? (Show your work) (2 pts)

Work and Answer:
Answer #1

15)

Paste the appropriate SPSS output. (3 pts)

T-Test

Group Statistics

Type of group

N

Mean

Std. Deviation

Std. Error Mean

Number of alphabets recall

Individual

10

15.8000

5.45283

1.72434

Group

10

14.3000

3.71334

1.17426

Independent Samples Test

Levene's Test for Equality of Variances

t-test for Equality of Means

F

Sig.

t

df

Sig. (2-tailed)

Mean Difference

Std. Error Difference

95% Confidence Interval of the Difference

Lower

Upper

Number of alphabets recall

Equal variances assumed

3.408

.081

.719

18

.481

1.50000

2.08620

-2.88294

5.88294

Equal variances not assumed

.719

15.870

.483

1.50000

2.08620

-2.92549

5.92549

16)

Paste appropriate SPSS graph.    (3 pts)

Required graph is given as below:

25.00 18 13 - 20.00 15.00 10.00 5.00 Group Individual Type of group

17)

Calculate the effect size given the SPSS output as shown in this week’s presentation. Show all of your work. (2 pts)

Effect size = (M1 – M2)/SD pooled = (15.8 – 14.3) / SD pooled

SD pooled = sqrt[(S1^2/n1)+(S2^2/n2)]

SD pooled = sqrt((5.45283^2/10)+(3.71334^2/10))

SD pooled = 2.086199

Effect size = (M1 – M2)/SD pooled = (15.8 - 14.3) /2.086199

Effect size = 0.719011

18)

Write an APA-style Results section based on your analyses. Remember to use complete sentences, include the statistical notation, and include a decision about the null hypothesis. If the results are significant, include the effect size and its interpretation as well. (4 pts)

For the given two sample independent samples t test for population means, the test statistic value is given as 0.719 with the p-value of 0.481. So, this P-value is greater than the alpha value 0.05, so we do not reject the null hypothesis. There is insufficient evidence to conclude that the average number of alphabets recall by two groups is different. The effect size is given as 0.7190 which indicate a medium effect.

19)

If the null hypothesis had been directional, what significance value would you report? (Show your work) (2 pts)

For the given two tailed t test, the p-value is given as 0.481, so the p-value for the directional or one tailed t test is given as 0.481/2 = 0.2405

Required p-value = 0.2405